OpenStudy (anonymous):

Find the value of 5x - 2y + 3z when x = 7, y = 11, and z = -1. Please show work

OpenStudy (texaschic101):

This is just a matter of subbing in known values and working it out. 5x - 2y + 3z.....when x = 7, y = 11, and z = -1 now we sub 5(7) - 2(11) + 3(-1) ca you finish ?

OpenStudy (texaschic101):

oops..typo...CAN you finish...lol

OpenStudy (wolf1728):

5x - 2y + 3z x=7 y=11 z =-1 5*7 -2*11 +3*-1 35 -22 -3 which equals 10

OpenStudy (anonymous):

35 - 22=13+-3= -10?

OpenStudy (texaschic101):

not quite.. 35 - 22 - 3 = 35 - 25 = 10

OpenStudy (anonymous):

how i dont get it

OpenStudy (texaschic101):

35 - 22 - 3 = 13 + (-3) = 13 - 3 = 10

OpenStudy (anonymous):

do 3 times -1

OpenStudy (anonymous):

it equalt negative 3

OpenStudy (anonymous):

nvm guys i get it now thx for the help

OpenStudy (texaschic101):

5(7) - 2(11) + 3(-1) = 35 - 22 - 3 --- now add your negatives 35 - 25 = 10